# Look before Transcription: End-to-End SlideASR with Visually-Anchored Policy Optimization

### v1

@@ 5 SlideASR-Bench: A Benchmark for entity-rich SlideASR Task
Our goal is to improve the accuracy of model transcription for domain-specific entities in the context of slides. Existing public SlideASR datasets, such as SlideSpeech Wang et al. (2024b) and ChineseLips Zhao et al. (2025), provide a valuable benchmark for real-world general-domain scenarios. However, we observe that these datasets often lack a sufficient density of domain-specific named entities in both speech and slides. This creates a significant challenge in training a model capable of handling entity-rich slides and speech, as well as objectively evaluating its performance in entity-rich SlideASR scenarios. To address this critical resource bottleneck, we constructed SlideASR-Bench, which includes two datasets: a synthetic dataset, SlideASR-S, for training and evaluation, and a challenging evaluation set, SlideASR-R, for assessing real-world performance. Table 2 presents the detailed information.
SlideASR-S. To train and evaluate models for entity-rich scenarios, we constructed SlideASR-S based on the ContextASR-Bench Wang et al. (2025) dataset. ContextASR-Bench leverages LLMs, such as DeepSeek-R1 Guo et al. (2025), to generate colloquial text rich in named entities based on seed text. The seed text is sourced from Named Entity Recognition (NER) datasets across multiple domains, such as medicine, culture, ecology. Then, models such as CosyVoice2 Du et al. (2024) and XTTS-v2 Casanova et al. (2024), which are text-to-speech models, are used to convert the generated text into natural and fluent speech.
We extract metadata for each sample from ContextASR-Bench, including the domain label $L_{\text{domain}}$ and the list of domain-specific entities $E=\{e_{1},e_{2},\dots,e_{k}\}$. Using the $L_{\text{domain}}$ and $E$ as input, we employ a LLM (e.g., Qwen2.5-14B-Instruct  [^1] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct) to generate a short descriptive text in slide style. The prompt guides the LLM to include a title and key points, ensuring that all entities from the original audio are naturally embedded in the generated text. The prompt can be found in Appendix  A. Finally, we use Python’s Matplotlib library to render the text into an image, simulating a presentation slide. Each generated image serves as synthetic visual context.
SlideASR-R. Although synthetic data is useful for model training and evaluation, there remains a certain domain gap between synthetic and real-world scenarios. To assess the model’s generalization ability in real, complex environments, we manually constructed a small-scale, high-quality, and challenging test set.
We collected 60 real presentation audio clips and corresponding slide images from publicly available academic report videos, covering four specialized domains: chemistry, medicine, biology, and artificial intelligence. For each sample, we manually annotated the data by carefully comparing the speech and slide image, identifying the domain-specific entities that appear in both. We named this dataset SlideASR-R (R for Real), which contains 200 domain-specific entities from real-world scenarios. Despite its relatively small size, it provides a highly challenging benchmark for assessing model performance in practical applications.
Table 2: Details of our proposed SlideASR-Bench.
#Sample #Entity #Hour
SlideASR-S (Train set) 6,413 44,240 67.3
SlideASR-S (Test set) 2,054 13,895 18.5
SlideASR-R 60 200 0.35
6 Experiment
6.1 Setup
@@
Figure 4: Attention visualization. Left: input image and transcribed audio text. Right: attention flows.
We fine-tune the Qwen2.5-Omni-3B and Qwen2.5-Omni-7B models on the SlideASR-S training set using the proposed VAPO algorithm. All experiments were run on 4 NVIDIA A100 GPUs with 80 GB of memory each. The weights of the reward functions, $\lambda_{1}$ to $\lambda_{4}$, are all set to 1.
We evaluate models on SlideSpeech Wang et al. (2024b) and SlideASR-Bench, using three settings: Contextless, where only audio is used; Slide text as context, a pipeline setting where OCR extracts text from slides and combines it with audio; and Slide image as context, an end-to-end setting that directly inputs both slide images and audio. We select mainstream LALMs Chu et al. (2024); Dinkel et al. (2025) and OLLMs Yao et al. (2024); Xu et al. (2025a, b) as baselines. Note that the same model may support multiple settings. For example, Qwen3-Omni-30B-A3B Xu et al. (2025b) can take audio alone as input or accept text or image as context. Following Zhao et al. (2025), we use PaddleOCR  [^2] https://github.com/PaddlePaddle/PaddleOCR. for OCR. See Appendix B for details of evaluation prompts, metrics and baseline models. Additionally, we present the results on ChineseLips Zhao et al. (2025), a real-world Chinese dataset, in Appendix C (Table 6) to demonstrate VAPO’s generalization in real-world scenarios.
6.2 Main Results
@@
Results on SlideSpeech. Table 3 shows the results on the real-world SlideSpeech dataset. It can be observed that baseline models generally perform poorly when incorporating slide information, with performance even degrading compared to the contextless setting. For instance, the Qwen3-Omni-30B-A3B Xu et al. (2025b), using either slide text or the slide image as context exhibits a higher WER compared to using audio alone. In contrast, our proposed VAPO method achieves the best results on the SlideSpeech dataset. The VAPO-7B model reaches a WER of 10.31 and a recall of 97.32 on the test set, outperforming baselines such as Qwen3-Omni-30B-A3B (Recall=95.5) and MaLa-ASR Yang et al. (2024) (WER=11.26).
Results on SlideASR-Bench. Table 4 presents the results on SlideASR-Bench. Unlike SlideSpeech, on SlideASR-Bench, all audio-only models struggle to recognize specialized named entities. For example, on the SlideASR-R dataset, even the strongest audio-only model, Qwen3-Omni-30B-A3B, reaches a NE-FNR of 41.09. This underscores the essential role of incorporating visual information in entity-rich ASR tasks. Additionally, baseline models fail to effectively utilize contextual information, whether in text or image form. For example, on SlideASR-S, when Qwen3-Omni-30B-A3B uses OCR text or slide images as context, although NE-WER decreases compared to audio-only input, the overall WER increases significantly. This aligns with our observations in Section 3, the model is not transcribing the speech properly but rather outputting the context content instead.
The VAPO method achieves significant improvements in recognizing key entities. For example, on the challenging real-world evaluation set SlideASR-R, VAPO-7B reduces the NE-FNR from the baseline best of 28.22 to 15.35. These results indicate that VAPO can accurately extract and anchor entity text from slides, thereby improving transcription accuracy. See Appendix D, for the successful cases and analysis of a failure case.
6.3 Ablation Results
Table 5: Ablation results on the SlideASR-R dataset.
ASR
Reward
OCR
Reward
VA
Reward
@@ 6.3 Ablation Results
Table 5 presents the ablation results for different reward functions. The results show that, the ASR reward enables the models to initially reference slide information. The OCR reward further improves the accuracy of slide text extraction, enhancing overall performance. Finally, adding the Visual Anchoring (VA) reward strengthens the models’ focus on key entities in <think>, achieving the best results. Additional ablation results for SlideSpeech (Table 7) and SlideASR-S (Table 8) are in Appendix E.1. Table 9 in Appendix E.2 presents the ablation results for the reward weights ($\lambda_{1}$$\lambda_{2}$$\lambda_{3}$$\lambda_{4}$), demonstrating the trade-off between overall transcription accuracy and entity recognition performance, and how different weight configurations impact this balance. The 1:1:1:1 configuration optimally balances both.
@@ 6.4 Attention Visualization
To examine the model’s behavior during final transcription, we visualize the attention weights of VAPO-7B on a SlideSpeech case, as shown in Fig 4. We observed that when the model generates speech transcription in <answer>, it refers to the entity information in <think>. For instance, for the key entity "Concentriq", after generating the token "Concent" <answer>, the model pays significant attention to the "ri" token in <think> and subsequently generates it. It then refers to the "q" token in the <think> block, enabling accurate and complete transcription of "Concentriq". A similar process occurs for the entity "proscia". This demonstrates that the model indeed references the slide information during transcription. Further cases can be found in Appendix F.
@@ 7 Conclusion
This paper identifies key failures in OLLMs when applied to the SlideASR task, where they often disregard audio input and function merely as OCR systems. To address this, we introduce Visually-Anchored Policy Optimization (VAPO), a novel training method that enforces a structured "Look before Transcription" reasoning process. By leveraging a <think><answer> format and multi-faceted reward functions, VAPO significantly enhances the model’s ability to integrate visual and auditory information. Furthermore, we developed SlideASR-Bench, an entity-rich benchmark, to facilitate more robust training and evaluation. Extensive experiments demonstrate that VAPO substantially improves the transcription accuracy, particularly for specialized terms, establishing a more effective end-to-end paradigm for SlideASR.
@@ 8 Limitations
Task Generalization. Our current approach is highly specialized for leveraging textual information from presentation slides and does not incorporate other visual cues, such as images of entities (e.g., pictures of specific drugs). In the future, we will adapt our "Look before Transcription" paradigm to handle more diverse multimodal environments, where various visual elements play a crucial role.
@@
Inference Efficiency: The structured reasoning process of VAPO introduces a computational overhead, resulting in higher inference latency compared to models of the same size (detailed in Appendix G). This makes VAPO most suitable for offline applications where accuracy is critical. However, this trade-off between latency and accuracy is often acceptable for offline transcription tasks where precision is paramount. We will explore strategies such as model distillation in our future research to improve efficiency, enabling its use in real-time applications.
@@ 9 Ethical Considerations
Appendix A Prompts for Generate SlideASR-S
The prompt for the LLM to generate text paragraphs based on a domain label and an entity list is as follows.
Prompt for Qwen2-14B-Instruct to generate slide text Given a domain label and a list of entities, generate a title and a paragraph for use in a PPT report, with the requirement that the paragraph includes these entities, Keep paragraphs within 150 words.
Domain label:
List of entities:
Output format:
Title
Paragraph

### v2

@@ 4 SlideASR-Bench: A Benchmark for Entity-rich SlideASR Task
Our primary goal is to enhance ASR accuracy for domain-specific entities within visual presentation contexts. While existing datasets like SlideSpeech Wang et al. (2024b) and ChineseLips Zhao et al. (2025), provide valuable benchmarks for general-domain scenarios, we observe that they often lack a sufficient density of domain-specific named entities. This scarcity creates a significant bottleneck for both training visually-anchored models and evaluating their capabilities in specialized scenarios. To bridge this gap, we introduce SlideASR-Bench, which comprises two distinct subsets: SlideASR-S, a large-scale synthetic corpus for training and evaluation, and SlideASR-R, a small real-world test set for stress-test. Detailed statistics are presented in Table 7.
SlideASR-S. To train and evaluate models for entity-rich scenarios, we constructed SlideASR-S by extending the ContextASR-Bench Wang et al. (2025) dataset. ContextASR-Bench leverages LLMs, such as DeepSeek-R1 Guo et al. (2025), to generate colloquial text rich in named entities based on seed text. The seed text is sourced from Named Entity Recognition (NER) datasets across multiple domains (e.g., medicine, culture, and ecology). Text-to-speech models Du et al. (2024) are then used to convert the generated text into natural and fluent speech.
We extract metadata for each sample from ContextASR-Bench, including the domain label $L_{\text{domain}}$ and the list of domain-specific entities $E$. Using the $L_{\text{domain}}$ and $E$ as input, we employ an LLM (e.g., Qwen2.5-14B-Instruct [^1] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct Team and others (2024)) to generate a short non-colloquial, formal text in slide style. The prompt guides the LLM to include a title and key points, ensuring that all entities from the original audio are naturally embedded in the generated text. Finally, we utilize Python’s Matplotlib library to render the generated text into slide images, generating a total of 8,467 samples (6,413 for training set, 2,054 for test set). We provide the prompt and data example in Appendix  A.
SlideASR-R. Furthermore, to assess the model’s generalization ability in real, complex environments, we manually constructed a small-scale, high-quality, and challenging test set. We collected 60 real presentation audio clips and corresponding slide images from publicly available academic report videos, covering four specialized domains: chemistry, medicine, biology, and artificial intelligence. For each sample, we manually annotated the data by carefully comparing the speech and slide image, identifying the domain-specific entities that appear in both. We named this dataset SlideASR-R, which contains 200 domain-specific entities from real-world scenarios.
5 Experiment
5.1 Experimental Setup
@@
[…]
Baselines & Settings. We benchmark mainstream LALMs: Qwen2-Audio Chu et al. (2024) and Mi-Dasheng Dinkel et al. (2025) and OLLMs: MiniCPM-o-2.6 Yao et al. (2024) Qwen2.5-Omni Xu et al. (2025a), Qwen3-Omni Xu et al. (2025b) on SlideSpeech Wang et al. (2024b) and SlideASR-Bench across three settings:
  1. 1.
    Contextless, which only use audio as inputs;
  2. 2.
    Slide text as context (Pipeline), employing PaddleOCR Cui et al. (2025) for textual context extraction following Zhao et al. (2025);
  3. 3.
    Slide image as context (End-to-end).
We provide more evaluation details in Appendix B. Additionally, we report results on the ChineseLips Zhao et al. (2025) dataset in Appendix C to demonstrate the generalization of VAPO.
Metrics. We utilize two sets of metrics consistent with prior works. For SlideSpeech Wang et al. (2024b), we employ four metrics: WER, which measures overall transcription; U-WER, the unbiased WER on non-keyword segments for general transcription quality; B-WER, the biased WER focusing on keyword spans; and Recall, which is the percentage of correctly recognized keywords. Furthermore, for SlideASR-Bench, consistent with ContextASR-Bench Wang et al. (2025), we focus on three metrics: WER; NE-WER, the WER of the named entity portion; and NE-FNR, the False Negative Ratio of named entities, which measures the proportion of missed ground-truth entities. See Appendix B.2 for specific calculation details of the metrics.
5.2 Main Results and Analysis
@@
Results on SlideSpeech. Table 2 presents the results on the real-world SlideSpeech dataset. Ideally, visual context should enhance recognition; however, we observe a detrimental effect in baseline models. Most baselines exhibit performance degradation when incorporating slide information compared to the contextless setting. For instance, the Qwen3-Omni Xu et al. (2025b) shows increased WER in both pipeline and end-to-end settings compared to using audio alone. In stark contrast, our VAPO method effectively leverages visual cues, achieving the best performance. VAPO-7B reaches a WER of 10.31 and a Recall of 97.32, significantly outperforming the Qwen3-Omni baseline and the previous SOTA MaLa-ASR Yang et al. (2024).
Results on SlideASR-Bench. Table 3 reports the performance on our benchmark. Two critical observations highlight the superiority of VAPO:
1) Audio-only models struggle with specialized entities. On the challenging SlideASR-R subset, even the strongest audio-only model (Qwen3-Omni) suffers from a high NE-FNR of 41.09, underscoring the necessity of visual presentation context for domain-specific terms.
2) Naive visual integration leads to catastrophic failure. On SlideASR-S, baseline OLLMs fail to utilize visual context effectively. Instead of improving, they suffer from severe visual interference. For example, Qwen3-Omni in the end-to-end setting yields an exploded WER of 101.45 and a worsened NE-WER of 59.64 (vs. 14.61 in audio-only), indicating that the model is hallucinating slide content rather than transcribing speech. Conversely, on SlideASR-S (en), VAPO-3B achieves a remarkable WER of 4.90. Furthermore, on the SlideASR-R, VAPO-7B reduces the NE-FNR from the best baseline’s 28.22 to 15.35. These results confirm that VAPO successfully enforces the “Look-then-Listen” paradigm, accurately anchoring visual entities without succumbing to interference. We present case study in Appendix D to provide visual comparison.
Table 4: Ablation results of rewards on SlideASR-R.
@@
Table 5: Sensitivity analysis of reward weights ($\lambda_{1}$$\lambda_{2}$$\lambda_{3}$$\lambda_{4}$) on SlideASR-S. The balanced 1:1:1:1 strategy achieves the best robustness across metrics (N-W: NE-WER, N-F: NE-FNR).
Weights SlideASR-S (en) SlideASR-S (zh)
WER N-W N-F WER N-W N-F
1:1:1:1 4.90 3.19 3.73 2.47 4.21 2.22
1:1:1:2 5.27 3.34 3.78 2.50 4.30 2.09
1:1:2:1 5.32 4.12 3.91 2.48 4.38 2.09
1:2:1:1 5.17 3.45 3.80 2.51 4.23 1.99
5.3 Ablation Study
Ablation on Reward Functions. Table 4 details the stepwise contributions on SlideASR-R. Compared to the baseline, introducing the ASR Reward yields the most significant boost, drastically reducing NE-WER by stabilizing generation. Adding the OCR Reward further refines the visual prior quality. Finally, incorporating the Visual Anchoring Reward achieves the best performance. This confirms that while ASR and OCR guarantees individual modality perception, the VA reward is essential for effective referencing. We provide additional ablation results in Appendix E.
Sensitivity to Reward Weights. To investigate hyperparameter sensitivity, we evaluate varying weight configurations on SlideASR-S, as detailed in Table 5. The results confirm that the balanced 1:1:1:1 scheme yields the most robust overall performance. We observe an inherent trade-off: doubling the VA reward ($\lambda_{4}$=2) enhances entity recall on the Chinese subset (improving NE-FNR to 2.09), but at the cost of increased overall WER due to over-aggressive anchoring. Conversely, double the ASR reward ($\lambda_{3}$=2) proves counterproductive, suppressing the model’s reliance on visual cues and significantly degrading NE-WER (e.g., rising from 3.19 to 4.12 on the English subset). Consequently, we adopt the equal weighting strategy for its simplicity and effectiveness.
5.4 Robustness against Mismatched Slides
To verify that VAPO performs valid visual anchoring rather than blindly copying slide content, we evaluated the model’s performance when fed with randomly mismatched slide images. As illustrated in Fig. 4, the model exhibits strong robustness: the overall WER remains stable at 6.70%, which is comparable to the audio-only baselines, including Qwen2.5-Omni-3B (6.42%) and the audio-only variant of VAPO-3B (7.40%). This indicates that VAPO effectively mitigates visual interference, safely falling back to auditory perception when visual context is unreliable.
Figure 4: Robustness analysis of VAPO-3B under mismatched visual context on SlideASR-S.
Table 6: Comparison between VAPO and SFT baselines on SlideASR-S. "SFT w/o think" is trained to directly predict the transcription, whereas "SFT w/ think" is supervised using the structured inference chain.
[…]
Qwen2.5-Omni-3B
+ SFT w/o think 8.44 8.18 9.30
+ SFT w/ think 6.60 6.47 7.10
+ VAPO 3.88 3.84 3.00
Qwen2.5-Omni-7B
+ SFT w/o think 10.58 10.51 12.77
+ SFT w/ think 6.73 5.56 6.53
+ VAPO 3.37 3.07 2.00
Figure 5: Attention visualization of VAPO-7B. The dense orange connections demonstrate the visual anchoring mechanism, showing that the model explicitly attends to extracted entities in <think> to guide the final transcription.
5.5 Impact of Training Paradigm: VAPO vs. Supervised Fine-Tuning (SFT)
To disentangle the contributions of the structured format from the reinforcement learning (RL) optimization, we compare three training settings on SlideASR-S: SFT w/o think (direct mapping), SFT w/ think (structured but SFT-only), and VAPO. Results in Table 6 highlights two key findings:
1) Efficacy of the Inference Chain. Introducing a structured inference chain yields significant gains, reducing the WER of Qwen2.5-Omni-7B from 10.58 to 6.73. This confirms that temporally decoupling the “Look” and “Listen” phases inherently mitigates visual interference.
2) Necessity of RL. VAPO further lowers the WER to 3.37. This indicates that while SFT teaches the format, it fails to establish a sufficient connection between <think> and <answer>. VAPO, driven by the multi-dimensional rewards, ensures the <think> block is effectively utilized for the final transcription.
@@ 5.6 Attention Visualization
To empirically validate the effectiveness of the visual anchoring mechanism, we visualize the attention weights of VAPO-7B on a sample from SlideSpeech, as shown in Fig. 5. The visualization reveals a distinct semantic reference pattern: during the generation of the <answer> block, the model heavily attends to the corresponding extracted text within the <think> block.
Specifically, take the proper noun “Concentriq” as an example. After generating the token “Concent” in <answer>, the model pays significant attention to the “ri” token in <think> and subsequently generates it. It then refers to the “q” token in the <think> block, enabling accurate and complete transcription of “Concentriq”. A similar process occurs for the entity “proscia”. This qualitative evidence confirms that VAPO successfully enforces the “Look-then-Listen” paradigm, where the model explicitly references the visual context prior. We provide more cases of attention visualization in Appendix F.
6 Related Works
Contextual ASR. The objective of contextual ASR is to incorporate contextual information, including domain labels, entity lists, and conversational history, into the speech recognition system in order to improve the recognition accuracy of domain-specific terminology Bai et al. (2024); Xiao et al. (2025); Zhou and Li (2025). Besides textual information, researchers have focused on leveraging visual information to enhance the performance of ASR models. For example, integrating lip movement information during the recognition process Ma et al. (2023); Rouditchenko et al. (2024); Shi et al. (2022). This study focuses on the SlideASR task Zhao et al. (2025); Wang et al. (2024b, a), which involves utilizing slide content as contextual information to support the model, given that slides in presentation scenarios generally contain information closely related to the spoken content. Most existing methods for SlideASR are based on the pipeline paradigm Wang et al. (2024a); Zhao et al. (2025); Wang et al. (2024b); Yu et al. (2024); Yang et al. (2024), which results in relatively complex systems. The objective of this study is to accomplish the task using an end-to-end approach.
Omni-modal Large Language Models. Recently, OLLMs Hurst et al. (2024); Fu et al. (2025); Yao et al. (2024); Xu et al. (2025a); Li et al. (2025a); Hu et al. (2025); Li et al. (2025b) have emerged, integrating vision, audio, and text by aligning their encoders during training for end-to-end processing. Models such as MiniCPM-o Yao et al. (2024) and Qwen2.5-Omni Xu et al. (2025a) have demonstrated strong multimodal performance. Benefiting from the unified modeling capability across visual and audio modalities, they are expected to solve the SlideASR task in an end-to-end manner. However, in practical scenarios, the models exhibit visual interference, for instance, they sometimes reproduce the textual content from the slides instead of generating the expected speech transcription.
Chain-of-Thought Reasoning. CoT reasoning is a breakthrough approach that enhances the reasoning capabilities of LLMs. Recent works Jaech et al. (2024); Guo et al. (2025) have shown that by using reinforcement learning algorithms Shao et al. (2024b); Rafailov et al. (2023); Schulman et al. (2017) to encourage models to generate intermediate reasoning steps before producing the final answer, performance on tasks involving arithmetic, commonsense, and symbolic reasoning can be significantly enhanced. This paradigm has also been extended to the multimodal domain Shao et al. (2024a); Xu et al. (2024); Ma et al. (2025); Lin et al. (2025); Diao et al. (2025), demonstrating the general effectiveness of making reasoning processes explicit.
@@ 7 Conclusion
In this work, we uncover a critical Visual Interference phenomenon in OLLMs applied to the SlideASR task, where models succumb to modality dominance and ignore auditory inputs. To address this, we introduce Visually-Anchored Policy Optimization (VAPO), a novel post-training method that enforces a human-like “Look-then-Listen” inference chain. By leveraging a structured <think><answer> format and multi-dimensional reward optimization, VAPO effectively decouples visual perception from auditory transcription. Furthermore, we contribute SlideASR-Bench, a comprehensive benchmark designed to address the data scarcity in entity-rich scenarios. Extensive experiments demonstrate that VAPO not only eliminates visual interference but also sets a new state-of-the-art performance, particularly in recognizing specialized domain entities.
@@ 8 Limitations
Task Generalization. Our current approach is highly specialized for leveraging textual information from presentation slides and does not incorporate other visual cues, such as images of entities (e.g., pictures of specific drugs). In the future, we will adapt our “Look-then-Listen” paradigm to handle more diverse multimodal environments, where various visual elements play a crucial role.
@@
Inference Efficiency. The structured reasoning process of VAPO introduces a computational overhead, resulting in higher inference latency compared to models of the same size (detailed in Appendix G). This makes VAPO most suitable for offline applications where accuracy is critical. However, this trade-off between latency and accuracy is often acceptable for offline transcription tasks where precision is paramount. We will explore strategies such as model distillation in our future research to improve efficiency, enabling its use in real-time applications.
@@ 9 Ethical Considerations
10 Acknowledgments
This work is supported by the National Key R&D Program of China (No. 2023YFC3310700), the Open Funding Programs of State Key Laboratory of AI Safety, the National Natural Science Foundation of China (No. 2576030), and the Beijing Nova Program under Grant 20250484899.
@@ 10 Acknowledgments
Appendix A Details of SlideASR-Bench
Statistics of SlideASR-Bench. Table 7 presents the key statistical metrics for SlideASR-Bench, including the number of samples, entities, and hours.
Table 7: Details of our proposed SlideASR-Bench.
Subset Sample Entity Hour
SlideASR-S (Training set) 6,413 44,240 67.3
SlideASR-S (Test set) 2,054 13,895 18.5
SlideASR-R 60 200 0.35
Prompt for Generating SlideASR-S. The prompt for the LLM to generate text paragraphs based on a domain label and an entity list is as follows. Fig. 6 shows an example of SlideASR-S.
Prompt for Qwen2-14B-Instruct to generate slide text Given a domain label and a list of entities, generate a title and a paragraph for use in a PPT report, with the requirement that the paragraph includes these entities, Keep paragraphs within 150 words. Domain label: {} List of entities: {} Output format: ### Title ### Paragraph
Figure 6: Data example of SlideASR-S.